OpenStudy (anonymous):

integral of sqrt(a^2-x^2)dx ... i got to the answer arcsin(x/a) + c ... but it appeared as wrong =\

OpenStudy (jamesj):

It is wrong because \[ \int \frac{1}{\sqrt{a^2 - x^2}} \ dx \ \ = \ \ \arcsin(x/a) + C \] For your integral, integrate by parts with du = 1 dx v = sqrt(a^2 - x^2) and take it from there.

OpenStudy (jamesj):

i.e., \[ \int \sqrt{a^2 - x^2} \ dx \ = \ x\sqrt{a^2 - x^2} - \int \frac{-x^2}{\sqrt{a^2-x^2}} \ dx \]

OpenStudy (anonymous):

You can solve it using trigonometric substitution. Substitute \(x=a\sin{u} \implies dx=a\cos{u}du\): \[\int\limits_{}^{}\sqrt{a^2-x^2}dx=\int\limits \sqrt{a^2-a^2\sin^2{u}} a \cos{u} du=a^2 \int\limits \cos^2u du.\] You can proceed from here.

OpenStudy (anonymous):

 You might need this triangle for your back substitution.